ELECTRON-SPIN AND TUNNELING EFFECTS IN THE MICROWAVE SPECTRUM OF SO_{2}–O_{2}

\underline{A.R. HIGHT WALKER}, G.T. FRASER, J.T. HOUGEN, C. LUGEZ, R.D. SUENRAM, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899; W. FAWZY, University of United Arab Emirates, Faculty of Science, P.O. Box 17551, Al-Ain, United Arab Emirates.

The rotational spectrum of the SO_{2}–O_{2} complex has been recorded between 6 GHz and 24 GHz using a pulsed-molecular-beam Fourier-transform microwave spectrometer. The spectrum is complicated by the spin of the triplet oxygen and tunneling of the two monomers within the complex. Approximately sixty a- and c-type transitions with J = 0 to 8 and K_{a} = 0,1,2. have been assigned and confirmed by combination differences. The observed transitions correlate to the lower (Ω = 0) energy component of the O_{2} spin-spin multiplet. Transitions associated with the higher energy (Ω = ±1) component have not been assigned, presumably due to the lower population of this state in the cold molecular beam (T_{r} =∼ 1 K). We note that in free O_{2} the spin-spin splitting is approximately 4 cm^{−1} (6 K). The ratio of the frequencies of the two observed ΔK = 1 subband origins is approximately 1.6, compared to a value of 3 expected for a rigid prolate top. A tunneling motion which reverses the sign of the c-type dipole moment component is used to explain this anomaly. Such a tunneling motion is anticipated from previous studies on Ar-SO_{2} and SO_{2} dimer. A fit of the observed transitions to a rigid rotor Hamiltonian with a tunneling term produces a standard deviation of 23 MHz and a tunneling splitting of 2.3 GHz. This standard deviation is significantly greater than the experimental precision of ∼1 kHz and is mainly attributed to the neglect of the electron spin. An alternative fit of this same data was carried out using a Hamiltonian which takes into account effects of electron spin and approximates a tunneling coefficient. This second fit produced errors on the order of 1 MHz and confirms that both the electron spin and the tunneling motion must be simultaneously considered. Future efforts are directed at developing a rotation-tunneling-spin Hamiltonian to model the spectrum. In addition, isotopic studies are being undertaken to determine the orientation of the SO_{2} and O_{2} subunits in the complex.